OpenStudy (anonymous):

Use the first derivative to determine where the function f(x)= -15x^2-10x+6 is increasing and decreasing. Use interval notion. Use U for a union (more than one interval) and leave it blank if the interval is the empty set.

OpenStudy (anonymous):

So, we have: \[ f(x)=-15x^2-10x+6 \]Thus: \[ f'(x)=-30x-10 \]So, we have points of inflection at: \[ f'(x)=-30x-10=0\implies\\ x=-\frac{1}{3} \]Using the second derivative test, we have: \[ f''(x)=-30<0 \]So it is a local maxima, which means that the function is increasing steadily until \(x=-\frac{1}{3}\), where it has an inflection point, and then steadily decreases. Et al.

OpenStudy (anonymous):

I love you ty ty ty

OpenStudy (anonymous):

Haha, sure thing.